OLIVER69:

I need some help with 12th grade finite math help... Questions 1-3

OLIVER69:

1 attachment
YourlocalRandom2000:

Oh i failed tht miserably

martai:

Well damm uhhh

strugglinginSchool27:

I have the answer for number one

strugglinginSchool27:

For this type of problem, it might be convenient to draw a table.

1 attachment
strugglinginSchool27:

Since w1 + w2 + w3 + w4 = 1, we have 0.5 + 3x + x + x = 1, and so 5x = 0.5, which gives x = 0.1. Therefore, w2 = 3x = 3(0.1) = 0.3